OpenStudy (anonymous):

true or false ? x^2+4/x+2 =x-2

OpenStudy (anonymous):

False

OpenStudy (lgbasallote):

\[\frac{x^2 - 4}{x+2} = x-2\] \[\frac{x^2 + 4}{x+2} \ne x-2\]

OpenStudy (anonymous):

 so